A field of length 58m and width 45m has four square corners of 8m. Find the perimeter of the excluding the corners
Answer by srinivas.g(540)

length = 58 m
effective length excluding corners of 8 m length = 58-8-8 ( each side consist of 2 8 m corners )
=42 m
effective width = 45-8-8
= 29 m
perimeter excluding 8 m corners = 2( effective length + effective width)
=2(42+29)
=2(71)
= 142 m
Result : perimeter = 142 m